OpenStudy (anonymous):

use laws of exponents to simplify. 5.6^7/4 / 5.6^4/5

OpenStudy (anonymous):

Is this the question ? \[\frac{5.6^\frac{7}{4}}{5.6^\frac{4}{5}}\]

OpenStudy (anonymous):

Assuming the above is what you mean , then what we need to do in the case of division, is subtract 5.6's exponents: \[=5.6^{(\frac{7}{4}-\frac{4}{5})}=5.6^{(\frac{35-16}{20})}=5.6^{(\frac{19}{20})}\]